A mass mA = 1.4kg , moving with velocity v? A=(5.0i
Answer #1

here ,

mA = 1.4 Kg

uA = 5i + 4.6 j - 3k m/s

mB = 3.8 Kg

vA = -3 + 3.4 k m/s

Now , using conservation of momentum ,

mA*uA = mA*vA + mB*vB

1.4 * (5i + 4.6 j - 3k) = 3.8 * vB + 1.4 * (-3i + 3.4 k)

solving for vB

vB = 2.95 i + 1.69 j -2.36 k m/s

the velocity of mass B is (2.95 i + 1.69 j -2.36 k) m/s